At what rate of interest, convertible semiannually, is $10,000 the present value of $1,000 paid at the end of every half-year for 6 years?
1
Expert's answer
2020-05-01T09:41:17-0400

Let the rate of interest be r % per annum.

"A=P(1+\frac r{2\times100})^{2n}"


"10000=1000(1+\frac r{200})^{12}"


"1+\frac{r}{200}=" "10^{1/12}\approx 1.211"

r"\approx" 42%.